OpenStudy (anonymous):

determine if the following ordered pairs are solutions to the linear equation 2x-y=3 A- (1,-1) B (4,6)

OpenStudy (nightwielder):

Plug it in and see which one works.

OpenStudy (anonymous):

idk how to

OpenStudy (nightwielder):

2(1)-(-1)=3 2+1=3 3=3 So (1,-1) works...

OpenStudy (nightwielder):

Now try the other one...

OpenStudy (nightwielder):

2(4)-6=3 8-6=3 \[2\neq3\]

OpenStudy (nightwielder):

So A would be your only solution.